a point on the rim of a 0.5 m radius fan blade has centripetal acceleration of 2.0 m/s^2. find the linear speed of a point 0.25 m from the center of the same wheel.
Answer #1

Centripetal acceleration=2 m/s²

Radius=0.5 m

Let angular velocity be w

0.5*w²=2

w=2 rad/s

Linear speed at radius of 0.25 m=0.25*2=0.5 m/s

This is thr required answer